Joey has a 97% chance of catching the early train to work each morning. What is the probability that he misses the train at least once in twenty mornings?
Answer by jim_thompson5910(35256)

Probability he catches the train 20 times in a row:
P(catches all 20) = (0.97)^20 = 0.54379

probability he misses at least one train:
P(misses at least one) = 1-P(catches all 20) = 1-0.54379 = 0.45621


Answer: 0.45621 or 45.621% (rough approximation)